the amount of strep in the plates is 100 ug/ml. if you have a 2% stock solution of streptomycin, how much would you need to add to prepare 1 litre of this medium?
Answer #1

Answer-

Stock solution of streptomycin = 2%(2g in 100 ml)

= 0.02g/ ml

=20mg/ml

C1  =20,000ug/ml

C2 = 100ug/ ml

V2= 1L = 1000ml

V1= ?

we can calculate amount of stock solution of streptomycin should be added in 1L of concentration 100ug/ml-

C1×V1=C2×V2

20,000×V1= 100× 1000

V1= (100×1000)÷20,000

=1,00,000÷20,000

V1=5ml

The 5ml of streptomycin stock solution is added to 995ml of this medium to prepare plates of 1000ml media with streptomycin concentration 100ug/ml.